A firm has 1,330,000 authorized shares, 510,000 issued shares, and 330,000 outstanding shares before completing a stock split. The shares have a par value of $0.04. The firm completes a reverse 2-to-1 split. How many shares does the firm have in treasury stock after the stock split?
Answer #1

Total Shares : 1,330,000

Issued shares : 510,000

Outstanidng No of shares : 330,000/- Nos

Treasury shares before split : 180,000 = (510,000- 330,000)

Shares issued in Split ratio of 2:1 : (330,000/2) = 165,000 Nos of shares

Tresury shares after split = 180,000 - 165,000 = 15,000 .

Number of shares in treasury stock after stock split :

Definition of treasury stock :